OpenStudy (anonymous):

[2.07] Solve 3|2x + 2| − 7 = 5 for x.

OpenStudy (anonymous):

3|2x + 2| − 7 = 5 l2x+2l = 4 \[(2x+2)^{2} = 16\] \[4x ^{2}+8x+4=16\]\[4x ^{2}+8x-12=0\]\[x ^{2}+2x-3=0\] (x - 1)(x + 3) = 0 therefore x=1 or x=-3

OpenStudy (anonymous):

thanks i think im getting it and just as i start to understand the next quesiton is [2.07] Create an absolute value equation that has two solutions. Explain why this equation has two solutions.

OpenStudy (anonymous):

what do i do now??!! :(

OpenStudy (sw050399):

3|2x + 2| - 7 = 5 3|2x + 2| = 12 |2x + 2| = 4 From here, you can get: 2x + 2 = 4 and 2x + 2 = -4 due to the absolute value bars Solving 2x + 2 =4: 2x + 2 = 4 2x = 2 x = 1 Solving 2x + 2 = -4: 2x = -6 x = -3

OpenStudy (anonymous):

yea that is correct

OpenStudy (anonymous):

infact the method u used is wat ur supposed to use, so dont use my working, ur working is the right one

OpenStudy (anonymous):

wow thank you

OpenStudy (anonymous):

:) keep it up